# Score identity Distillation: Exponentially Fast Distillation of Pretrained Diffusion Models for One-Step Generation

### v2

@@ 4.1 Model-based Explicit Score Matching (MESM)
We refer to $\delta_{\phi,\theta}(\boldsymbol{x}_{t})$ as score difference and designate its L2 norm ${\mathcal{L}}_{\theta}$ as the model-based explicit score-matching (MESM) loss, also known in the literature as a Fisher divergence (Lyu, 2009; Holmes & Walker, 2017; Yang et al., 2019; Yu & Zhang, 2023). This differs from the ESM loss defined in (6) as the expectation is computed with respect to the diffused fake data distribution $p_{\theta}(\boldsymbol{x}_{t})$ rather than diffused real data distribution $p_{\text{data}}(\boldsymbol{x}_{t})$.
@@ 4.3 Loss Approximation via Projected Score Matching
${\mathcal{L}}_{\theta}=\mathbb{E}_{q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g},t)p_{\theta}(\boldsymbol{x}_{g})}\left[\sigma_{t}^{-2}{\delta_{\phi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})^{T}(f_{\phi}(\boldsymbol{x}_{t},t)-\boldsymbol{x}_{g})}\right]$.
@@ 6 Conclusion
Acknowledgments
The authors would like to thank Dr. Zhuoran Yang and Weijian Luo for their valuable comments and suggestions.
@@ Appendix A Ablation Study and Parameter Settings
Setting of $\beta_{1}$. We investigate the $\beta_{1}$ parameter of the Adam optimizer for the generator score network $f_{\psi}$ and the generator $G_{\theta}$ by setting it as either $\beta_{1}=0$, the value used in StyleGAN2 (Karras et al., 2020) and Diff-Instruct (Luo et al., 2023c), or $\beta_{1}=0.9$, a commonly used value. We find that setting $\beta_{1}=0.9$ for $f_{\psi}$ often does not result in convergence, so we retain $\beta_{1}=0$ for $f_{\psi}$ for all datasets. For learning $G_{\theta}$, we did not observe significant benefits between setting $\beta_{1}=0$ and $\beta_{1}=0.9$, except for the FFHQ dataset, where the FID improved by more than 0.15 when changing from $\beta_{1}=0$ to $\beta_{1}=1$. Therefore, we set $\beta_{1}=1$ for FFHQ while retaining $\beta_{1}=0$ for all other datasets.
@@
Figure 3: Ablation Study of $\alpha$: Each plot illustrates the relation between the performance, measured by Inception Score and FID vs. the number of every thousand images processed during the distillation of the EDM model pretrained on CIFAR-10 (unconditional), across varying values of $\alpha$. The study underscores the impact of $\alpha$ on both training efficiency and generative fidelity, leading us to select $\alpha\in\{1.0,1.2\}$ for all subsequent experiments.
@@ Appendix B Algorithm Box
   Sample $\boldsymbol{z}\sim\mathcal{N}(0,\mathbf{I})$ and $\boldsymbol{x}_{g}=G_{\theta}(\sigma_{\text{init}}\boldsymbol{z})$
   Sample $t\sim p(t),~\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(0,\mathbf{I})$ and $\boldsymbol{x}_{t}=\boldsymbol{x}_{g}+\sigma_{t}\boldsymbol{\epsilon}_{t}$
   Update $\psi$ with Equation 10:
       $\mathcal{L}_{\psi}={\gamma(t)}\|f_{\psi}(\boldsymbol{x}_{t},t)-\boldsymbol{x}_{g}\|_{2}^{2}$
       $\psi=\psi-\eta\nabla_{\psi}\mathcal{L}_{\psi}$
@@
   Sample $t\sim\mbox{Unif}[0,t_{\max}/1000]$, compute $\sigma_{t}$ with Equation 26, compute $\omega_{t}$ with Equation 25, and update $G_{\theta}$ with Equation 23:
@@ Appendix C Proofs
$\nabla_{\boldsymbol{x}_{t}}\ln p_{\theta}(\boldsymbol{x}_{t})$ $=\frac{\int\nabla_{\boldsymbol{x}_{t}}q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g})p_{{\theta}}(\boldsymbol{x}_{g})d\boldsymbol{x}_{g}}{p_{\theta}(\boldsymbol{x}_{t})}$
@@
$\nabla_{\boldsymbol{x}_{t}}\ln p_{\text{data}}(\boldsymbol{x}_{t})$ $=\frac{\int\nabla_{\boldsymbol{x}_{t}}q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{0})p_{{\phi}}(\boldsymbol{x}_{0})\mathrm{d}\boldsymbol{x}_{0}}{p_{\phi}(\boldsymbol{x}_{t})}$ (28)
@@
$=\mathbb{E}_{\boldsymbol{x}_{t}\sim p_{\theta}(\boldsymbol{x}_{t})}[u^{T}(\boldsymbol{x}_{t})\frac{\nabla_{\boldsymbol{x}_{t}}p_{\theta}(\boldsymbol{x}_{t})}{p_{\theta}(\boldsymbol{x}_{t})}]$
@@
$=\int u^{T}(\boldsymbol{x}_{t})\int\nabla_{\boldsymbol{x}_{t}}q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g})p_{\theta}(\boldsymbol{x}_{g})\mathrm{\mathrm{}}d\boldsymbol{x}_{g}d\boldsymbol{x}_{t}$
$=\int u^{T}(\boldsymbol{x}_{t})\int q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g})\nabla_{\boldsymbol{x}_{t}}\ln q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g})p_{\theta}(\boldsymbol{x}_{g})\mathrm{d}\boldsymbol{x}_{g}d\boldsymbol{x}_{t}$
@@
① $=\frac{1}{\sigma_{t}^{2}}\mathbb{E}_{\boldsymbol{x}_{t}\sim p_{\theta}(\boldsymbol{x}_{t})}[\delta_{\phi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})^{T}(\mathbb{E}[\boldsymbol{x}_{0}\,|\,\boldsymbol{x}_{t}]-\boldsymbol{x}_{t})]$
@@
② $=\mathbb{E}_{\boldsymbol{x}_{g}\sim p_{\theta}(\boldsymbol{x}_{g})}\mathbb{E}_{\boldsymbol{x}_{t}\sim q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g},t)}[\delta_{\phi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})^{T}\nabla_{\boldsymbol{x}_{t}}\ln p_{\theta}(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g}))]$
@@ Appendix D Analytic study of the toy example
We prove the conclusions in Propositions 4 and 6. Given $p_{\text{data}}(\boldsymbol{x}_{0})=\mathcal{N}(\mathbf{0},\mathbf{I})$, $p_{\text{data}}(\boldsymbol{x}_{t})=\mathcal{N}(0,1+\sigma_{t}^{2})$, and $q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g})=\mathcal{N}(\boldsymbol{x}_{g},\sigma_{t}^{2}),~p_{\theta}(\boldsymbol{x}_{g})=\mathcal{N}(\theta,1)$, the optimal value of $\theta$ would be $\theta^{*}=0$. The score
@@
Hence, the difference between the scores is $\delta_{\phi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})=-\frac{\theta}{1+\sigma_{t}^{2}}$. By Tweedie’s formula
@@
which means $\psi^{*}=\theta$, then by Equation 12 we have
@@
Figure 7: Uncoditional CIFAR-10 random images generated with SiD (FID: 1.923).
Figure 8: Label conditioning CIFAR-10 random images generated with SiD (FID: 1.710)
@@
Figure 10: FFHQ random images generated with SiD (FID: 1.550)
Figure 11: AFHQ random images generated with SiD (FID: 1.628)

### v3

@@ 4.1 Model-based Explicit Score Matching (MESM)
We refer to $\delta_{\phi,\theta}(\boldsymbol{x}_{t})$ as score difference and designate its expected L2 norm ${\mathcal{L}}_{\theta}$ as the model-based explicit score-matching (MESM) loss, also known in the literature as a Fisher divergence (Lyu, 2009; Holmes & Walker, 2017; Yang et al., 2019; Yu & Zhang, 2023). This differs from the ESM loss defined in (6) as the expectation is computed with respect to the diffused fake data distribution $p_{\theta}(\boldsymbol{x}_{t})$ rather than diffused real data distribution $p_{\text{data}}(\boldsymbol{x}_{t})$.
@@ 4.3 Loss Approximation via Projected Score Matching
[…]
@@ 6 Conclusion
Acknowledgments
The authors would like to thank Dr. Zhuoran Yang and Weijian Luo for their valuable comments and suggestions. M. Zhou, H. Zheng, and Z. Wang acknowledge the support of NSF-IIS 2212418 and NIH-R37 CA271186.
@@ Appendix A Ablation Study and Parameter Settings
Setting of $\beta_{1}$. We investigate the $\beta_{1}$ parameter of the Adam optimizer for the generator score network $f_{\psi}$ and the generator $G_{\theta}$ by setting it as either $\beta_{1}=0$, the value used in StyleGAN2 (Karras et al., 2020) and Diff-Instruct (Luo et al., 2023c), or $\beta_{1}=0.9$, a commonly used value. We find that setting $\beta_{1}=0.9$ for $f_{\psi}$ often does not result in convergence, so we retain $\beta_{1}=0$ for $f_{\psi}$ for all datasets. For learning $G_{\theta}$, we did not observe significant benefits between setting $\beta_{1}=0$ and $\beta_{1}=0.9$, except for the FFHQ dataset, where the FID improved by more than 0.15 when changing from $\beta_{1}=0$ to $\beta_{1}=0.9$. Therefore, we set $\beta_{1}=0.9$ for $G_{\theta}$ in FFHQ while retaining $\beta_{1}=0$ for all other datasets.
@@
Figure 3: Ablation Study of $\alpha$: Each plot illustrates the relation between the performance, measured by Inception Score and FID vs. the number of training iterations during the distillation of the EDM model pretrained on CIFAR-10 (unconditional), across varying values of $\alpha$. The study underscores the impact of $\alpha$ on both training efficiency and generative fidelity, leading us to select $\alpha\in\{1.0,1.2\}$ for all subsequent experiments.
@@ Appendix B Algorithm Box
   Sample $\boldsymbol{z}\sim\mathcal{N}(0,\mathbf{I})$ and let $\boldsymbol{x}_{g}=G_{\theta}(\sigma_{\text{init}}\boldsymbol{z})$; Sample $t\sim p(t)$ and $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(0,\mathbf{I})$, and let $\boldsymbol{x}_{t}=\boldsymbol{x}_{g}+\sigma_{t}\boldsymbol{\epsilon}_{t}$; Update $\psi$ with Equation 10:
       $\hat{\mathcal{L}}_{\psi}={\gamma(t)}\|f_{\psi}(\boldsymbol{x}_{t},t)-\boldsymbol{x}_{g}\|_{2}^{2}$
       $\psi=\psi-\eta\nabla_{\psi}\hat{\mathcal{L}}_{\psi}$
@@
   Sample $\boldsymbol{z}\sim\mathcal{N}(0,\mathbf{I})$ and let $\boldsymbol{x}_{g}=G_{\theta}(\sigma_{\text{init}}\boldsymbol{z})$; Sample $t\sim\mbox{Unif}[0,t_{\max}/1000]$, compute $\sigma_{t}$ with Equation 26, compute $\omega_{t}$ with Equation 25, and let $\boldsymbol{x}_{t}=\boldsymbol{x}_{g}+\sigma_{t}\boldsymbol{\epsilon}_{t}$; Update $G_{\theta}$ with Equation 23:
@@ Appendix C Proofs
$\nabla_{\boldsymbol{x}_{t}}\ln p_{\theta}(\boldsymbol{x}_{t})$ $=\frac{\int\nabla_{\boldsymbol{x}_{t}}q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g})p_{{\theta}}(\boldsymbol{x}_{g})\mathrm{d}\boldsymbol{x}_{g}}{p_{\theta}(\boldsymbol{x}_{t})}$
@@
$=\mathbb{E}_{\boldsymbol{x}_{t}\sim p_{\theta}(\boldsymbol{x}_{t})}\left[u^{T}(\boldsymbol{x}_{t})\frac{\nabla_{\boldsymbol{x}_{t}}p_{\theta}(\boldsymbol{x}_{t})}{p_{\theta}(\boldsymbol{x}_{t})}\right]$
@@
$=\int u^{T}(\boldsymbol{x}_{t})\int\nabla_{\boldsymbol{x}_{t}}q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g})p_{\theta}(\boldsymbol{x}_{g})\mathrm{d}\boldsymbol{x}_{g}\mathrm{d}\boldsymbol{x}_{t}$
$=\int u^{T}(\boldsymbol{x}_{t})\int q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g})\nabla_{\boldsymbol{x}_{t}}\ln q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g})p_{\theta}(\boldsymbol{x}_{g})\mathrm{d}\boldsymbol{x}_{g}\mathrm{d}\boldsymbol{x}_{t}$
@@
① $=\frac{1}{\sigma_{t}^{2}}\mathbb{E}_{\boldsymbol{x}_{t}\sim p_{\theta}(\boldsymbol{x}_{t})}\left[\delta_{\phi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})^{T}(\mathbb{E}[\boldsymbol{x}_{0}\,|\,\boldsymbol{x}_{t}]-\boldsymbol{x}_{t})\right]$
@@
② $=\mathbb{E}_{\boldsymbol{x}_{g}\sim p_{\theta}(\boldsymbol{x}_{g})}\mathbb{E}_{\boldsymbol{x}_{t}\sim q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g},t)}[\delta_{\phi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})^{T}\nabla_{\boldsymbol{x}_{t}}\ln q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g}))]$
@@ Appendix D Analytic study of the toy example
We prove the conclusions in Propositions 4 and 6. Given $p_{\text{data}}(\boldsymbol{x}_{0})=\mathcal{N}(\mathbf{0},\mathbf{I})$, $p_{\theta}(\boldsymbol{x}_{g})=\mathcal{N}(\theta,\mathbf{I})$, $q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{0})=\mathcal{N}(\boldsymbol{x}_{t};\boldsymbol{x}_{0},\sigma_{t}^{2}\mathbf{I})$, and $q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g})=\mathcal{N}(\boldsymbol{x}_{t};\boldsymbol{x}_{g},\sigma_{t}^{2}\mathbf{I})$, we have $p_{\text{data}}(\boldsymbol{x}_{t})=\mathcal{N}(0,(1+\sigma_{t}^{2})\mathbf{I})$ and $p_{\theta}(\boldsymbol{x}_{t})=\mathcal{N}(\theta,(1+\sigma_{t}^{2})\mathbf{I})$. The optimal value of $\theta$ would be $\theta^{*}=0$. The score can be expressed as
@@
Hence, the difference between the scores is $\delta_{\phi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})=-\frac{\theta}{1+\sigma_{t}^{2}}$. By applying Tweedie’s formula as described in Identities 1 and 2, we obtain
@@
which means $\psi^{*}(\theta)=\theta$, then by Equation 12 we have
@@
Figure 7: Uncoditional CIFAR-10 32X32 random images generated with SiD (FID: 1.923).
Figure 8: Label conditioning CIFAR-10 32X32 random images generated with SiD (FID: 1.710)
@@
Figure 10: FFHQ 64X64 random images generated with SiD (FID: 1.550)
Figure 11: AFHQ-V2 64X64 random images generated with SiD (FID: 1.628)